Reports 1-1 of 1 Clear search Modify search
Virgo Runs (ER16)
berni - 15:01 Tuesday 02 April 2024 (63811) Print this report
Operator Report - Morning shift

ITF found in Science Mode; at 7:00 UTC I set Maintenance.

Below the list of the activities (known in control room):

  • cleaning operations of experimental areas;
  • cryotrap refill
  • TCS
    • at 6:30 UTC I performed the checks of TCS powers (see attachment);
    • at 8:00 UTC I acquired the thermal camera reference;
    • chiller refill by Julien;
  • DET
    • 6:54 UTC I performed the lock of the OMC in single bounce;
    • 7:04 UTC I performed the scan of the OMC in single bounce;
  • ISYS: at 7:35 UTC I performed the check of the IMC working point; the DMS showed a red flag. Experts have been informed;
  • OSD: at 8:35 UTC, under request of Fiodor, I performed the measurement of the cold RTL;
  • PAY: at 9:35 UTC swap of the NI power supply of the PSD (NI LC manually opened); completed at 10:00 UTC;


At 10:00 UTC all the activities were concluded and we started to relock; the ISC team tool the occasion to verify the lock acquisition.

At around 10:44 UTC Valerio made some tests and cheks on the PR.

Unfurtunately at the end of the shift the wind increased a lot preventing the lock, relock in progress.


Images attached to this report
Non-image files attached to this report
Comments to this report:
gouaty - 9:05 Wednesday 03 April 2024 (63822) Print this report

The attached figure shows the lock of the OMC and the OMC scan in single bounce performed yesterday.

The OMC was strongly misaligned. The "0.0" offsets for the B5_QD2 signals are clearly not good to have the OMC aligned in single bounce.

Images attached to this comment
Search Help